OpenStudy (roxy.girl):

Solve 4.2m + 4 = 25 @Qwertty123

rishavraj (rishavraj):

wht can u do to isolate the variable term??? @Roxy.girl

OpenStudy (qwertty123):

^ what can you do?

OpenStudy (roxy.girl):

you add -4?

OpenStudy (qwertty123):

Okay Reorder the terms: 4 + 4.2m = 25 Solving 4 + 4.2m = 25 Move all terms containing m to the left, all other terms to the right.

rishavraj (rishavraj):

so u will be having 4.2 m = 21 @Roxy.girl now wht can u do??

rishavraj (rishavraj):

now addition or subtraction isn't gonna help.. think differently @Roxy.girl u nees to get rid of 4.2

OpenStudy (roxy.girl):

you would divide. 4.2?

OpenStudy (qwertty123):

Yes! Add '-4' to each side of the equation. 4 + -4 + 4.2m = 25 + -4 Combine like terms: 4 + -4 = 0 0 + 4.2m = 25 + -4 4.2m = 25 + -4 Combine like terms: 25 + -4 = 21 4.2m = 21 Divide each side by '4.2'.

OpenStudy (qwertty123):

lol and yes XD

rishavraj (rishavraj):

@Roxy.girl good work and yeah

OpenStudy (qwertty123):

So what you get when you divide?

OpenStudy (roxy.girl):

5?

rishavraj (rishavraj):

@Roxy.girl was jst wondering do u give any try to questions before posting it on here??? no hard feelings :))

OpenStudy (qwertty123):

yes! m = 5 Nice job Meghan!

OpenStudy (roxy.girl):

yes I did... Qwertty is my tutor. XD but she is in a different room then where I am ATM.

OpenStudy (roxy.girl):

I am also studying for a test coming up.

rishavraj (rishavraj):

ooo cool xD

OpenStudy (roxy.girl):

lol yes we know each other in real life.

OpenStudy (qwertty123):

yup :) How about on more before we go on to another subject :)

rishavraj (rishavraj):

and I thought Qwertty is a guy :P

OpenStudy (qwertty123):

Every one does ^.^ So I get use to it.. lol

OpenStudy (roxy.girl):

lol xD Everyone does.